From the data in the table, calculate the final [FeSCN2+]:

Beaker

0.200 M Fe(NO3)3
(mL)

0.0020 M SCN
(mL)

H2O
(mL)

[FeSCN2+]

1

5.0

4.0

41.0

2

5.0

3.0

42.0

3

5.0

2.0

43.0

4

5.0

1.0

44.0
Answer #1

The reaction between Fe(NO3)3 and SCN to form FeSCN2+ is

Fe(NO3)3 + SCN -------> FeSCN2+

That means here all the components are euqimolar.i.e if we take 1M of Fe(NO3)3 and 1M of SCN , we will get 1 M of FeSCN2+

Now given

1- Beaker 1-

Initial Fe(NO3)3 taken = M1 = 0.200 M and V1 = 5.0 ml

Initial SCN taken = M1 = 0.0020 M and V1 = 4.0 ml

Water added = 41 ml

Now after the addition of these 3, the final volume = V2 = 5.0 ml +  4.0 ml + 41 ml = 50 ml

Thus the new concentration of Fe(NO3)3 in the total solution = M2 = M1V1/V2 = 0.200 M*5.0 ml / 50 ml = 0.02‬ M

Similarly new concentration of SCN  in the total solution = M2 = M1V1/V2 = 0.0020 M*4.0 ml / 50 ml = 0.00016‬ M

Thus in the final solution we have

0.02‬ M Fe(NO3)3 and 0.00016‬ M SCN . Since here SCN is present in less amount thus SCN is the limiting reagant.

Hence 0.00016‬ M SCN will react with 0.00016‬ M Fe(NO3)3 to form   0.00016‬ M FeSCN2+. The extra Fe(NO3)3 will remain as such in solution unreacted.

Thus [FeSCN2+] =  0.00016‬ M

2- Similarly-

Beaker 2-

Initial Fe(NO3)3 taken = M1 = 0.200 M and V1 = 5.0 ml

Initial SCN taken = M1 = 0.0020 M and V1 = 3.0 ml

Water added = 42 ml

Now after the addition of these 3, the final volume = V2 = 5.0 ml +  3.0 ml + 43 ml = 50 ml

Thus the new concentration of Fe(NO3)3 in the total solution = M2 = M1V1/V2 = 0.200 M*5.0 ml / 50 ml = 0.02‬ M

Similarly new concentration of SCN  in the total solution = M2 = M1V1/V2 = 0.0020 M*3.0 ml / 50 ml = 0.00012 M

Thus in the final solution we have

0.02‬ M Fe(NO3)3 and 0.00012‬ M SCN . Since here SCN is present in less amount thus SCN is the limiting reagant.

Hence 0.00012‬ M SCN will react with 0.00012‬ M Fe(NO3)3 to form   0.00012‬ M FeSCN2+. The extra Fe(NO3)3 will remain as such in solution unreacted.

Thus [FeSCN2+] =  0.00012 M

3-

Beaker 3-

Initial Fe(NO3)3 taken = M1 = 0.200 M and V1 = 5.0 ml

Initial SCN taken = M1 = 0.0020 M and V1 = 2.0 ml

Water added = 43 ml

Now after the addition of these 3, the final volume = V2 = 5.0 ml +  2.0 ml + 43 ml = 50 ml

Thus the new concentration of Fe(NO3)3 in the total solution = M2 = M1V1/V2 = 0.200 M*5.0 ml / 50 ml = 0.02‬ M

Similarly new concentration of SCN  in the total solution = M2 = M1V1/V2 = 0.0020 M*2.0 ml / 50 ml = 0.00008‬‬ M

Thus in the final solution we have

0.02‬ M Fe(NO3)3 and 0.00008‬‬ M SCN . Since here SCN is present in less amount thus SCN is the limiting reagant.

Hence 0.00008‬‬ M SCN will react with 0.00008‬‬ M Fe(NO3)3 to form   0.00008‬‬ M FeSCN2+. The extra Fe(NO3)3 will remain as such in solution unreacted.

Thus [FeSCN2+] =  0.00008‬‬ M

4-

Beaker 4-

Initial Fe(NO3)3 taken = M1 = 0.200 M and V1 = 5.0 ml

Initial SCN taken = M1 = 0.0020 M and V1 = 1.0 ml

Water added = 44 ml

Now after the addition of these 3, the final volume = V2 = 5.0 ml +  1.0 ml + 44 ml = 50 ml

Thus the new concentration of Fe(NO3)3 in the total solution = M2 = M1V1/V2 = 0.200 M*5.0 ml / 50 ml = 0.02‬ M

Similarly new concentration of SCN  in the total solution = M2 = M1V1/V2 = 0.0020 M*1.0 ml / 50 ml = 0.00004‬‬ M

Thus in the final solution we have

0.02‬ M Fe(NO3)3 and 0.00004‬‬ M SCN . Since here SCN is present in less amount thus SCN is the limiting reagant.

Hence 0.00004‬ M SCN will react with 0.00004‬‬ M Fe(NO3)3 to form   0.00004‬ M FeSCN2+. The extra Fe(NO3)3 will remain as such in solution unreacted.

Thus [FeSCN2+] =  0.00004‬ M